Please help me with this problem. I have the answer sheet but I am stuck on some of the steps and can't seem to get to the final answer.
5/(x^2-x)-1/(2x-2)=1/x

Answer by richwmiller(17219)

Too bad you don't share the answer you have !
If you share yours I might sh
are mine!
5/(x^2-x)-1/(2x-2)=1/x
5/x*(x-1)-1/2*(x-1)=1/x
(10-x)/2x(x-1) = 1/x
(x-4)/((x-1)*x) = 0
x-4=0
x=4